# Diffusion Posterior Sampling for Nonlinear CT Reconstruction

### v1

@@ Abstract
Diffusion models have been demonstrated as powerful deep learning tools for image generation in CT reconstruction and restoration. Recently, diffusion posterior sampling, where a score-based diffusion prior is combined with a likelihood model, has been used to produce high quality CT images given low-quality measurements. This technique is attractive since it permits a one-time, unsupervised training of a CT prior; which can then be incorporated with an arbitrary data model. However, current methods only rely on a linear model of x-ray CT physics to reconstruct or restore images. While it is common to linearize the transmission tomography reconstruction problem, this is an approximation to the true and inherently nonlinear forward model. We propose a new method that solves the inverse problem of nonlinear CT image reconstruction via diffusion posterior sampling. We implement a traditional unconditional diffusion model by training a prior score function estimator, and apply Bayes rule to combine this prior with a measurement likelihood score function derived from the nonlinear physical model to arrive at a posterior score function that can be used to sample the reverse-time diffusion process. This plug-and-play method allows incorporation of a diffusion-based prior with generalized nonlinear CT image reconstruction into multiple CT system designs with different forward models, without the need for any additional training. We develop the algorithm that performs this reconstruction, including an ordered-subsets variant for accelerated processing and demonstrate the technique in both fully sampled low dose data and sparse-view geometries using a single unsupervised training of the prior.
@@ 1 INTRODUCTION
Neural networks have been widely investigated for CT reconstruction and have shown impressive performance capabilities [1, 2, 3]. Recently, diffusion models have been shown to be particularly powerful deep learning tools for CT reconstruction and restoration [4, 5, 6, 7, 8]. Such methods have generally been based on denoising diffusion probabilistic models (DDPM) [9] and score-based generative diffusion models through stochastic differential equations (SDEs) [10]. Most diffusion models for CT reconstruction have been trained in a supervised manner, where the conditional data input is known in both training and generation [7, 8, 11]. Recent research on image denoising and restoration has demonstrated the potential of unsupervised training by leveraging the power of posterior sampling, where the conditional data input is only available during the reverse-time diffusion process [12, 13, 5]. The unsupervised training does not assume a fixed measurement process during training, and can thus be flexibly incorporated with different measurement models without retraining. In these efforts, the measurement model has been presumed to be linear. Lopez-Montes et al.[14] have recently applied such an approach using a linearized model to tomographic data.
@@
It is well-known that the transmission tomography forward model that relates the underlying attenuation values to the measurements is nonlinear. While the model is commonly linearized for mathematical convenience, such approximation can result in deviation from the true underlying data model and can cause problems in many circumstances including models for very low dose data [15], beam-hardening [16], and detector blur [17]. In this work, we propose an approach for nonlinear CT image reduction using diffusion posterior sampling (which we will refer to as DPS Nonlinear). Specifically, we seek to solve the inverse problem of CT image reconstruction via DPS with a score-based diffusion prior and a physics-driven measurement likelihood term. This method guides the updates of the reverse-time diffusion process so that the final results are consistent with the raw (nonlinear) measurements from a CT imaging system. Our approach trains a traditional score-based diffusion model for unconditional CT image generation and applies Bayes rule by adding the gradient of the log-likelihood (i.e.) from the non-linear physical model to achieve the posterior score function necessary for DPS. The method is plug-and-play, meaning it does not require any extra training for application to a different CT system measurement model.
2 METHODS
2.1 Diffusion Posterior Sampling using a Nonlinear Physical Model
@@
The proposed DPS Nonlinear method combines diffusion posterior sampling with a nonlinear physical model. This approach requires estimation of the prior distribution of CT images using a score-based generative model [9, 10]. Specifically, a diffusion model is adopted wherein a forward stochastic process is defined where an image is perturbed by successively added noise. A reverse process is then defined where we may sample images from the estimated prior distribution starting from a noise image. Specifically, we follow prior work [9, 10] where the forward diffusion process is governed by
$\mathbf{x_{t}}=\sqrt{\bar{\alpha}_{t}}\mathbf{x_{0}}+(1-\bar{\alpha}_{t})\mathbf{\epsilon_{noise}}\;\;\;t\in(0,T],$ (1)
@@
where $\mathbf{\epsilon_{noise}}\sim N(0,I)$ is Gaussian noise, $\mathbf{x_{0}}\sim\mathbf{p_{data}}$ denote a true image sampled from a training dataset, $\bar{\alpha}_{t}$ is a variance scheduler such that the final noise is close to a predefined noise distribution $\mathbf{\pi(x)}$, and the process is defined on $t\in(0,T]$. The final sample $\mathbf{p_{t=T}}$ usually follows the predefined noise distribution $\mathbf{\pi(x)}$ and it does not contain any information of training dataset $\mathbf{p_{data}}$ or $\mathbf{p_{t=0}}$. To recover the data in a tractable distribution, which can be achieved by reversing the perturbation forward process in Equation (1) to get the reverse-time SDE given by
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}})]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T],$ (2)
@@
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}}|\mathbf{y})]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (3)
@@
$\mathbf{d\mathbf{x_{t}}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}})+\nabla_{\mathbf{x_{t}}}\log p(\mathbf{y}|\mathbf{x_{t}}))]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (4)
@@
Following previous score matching methods[10, 18], the prior score $\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}})$ is estimated by a neural network to form a score model $\mathbf{s_{\theta}}(\mathbf{x_{t}},t)$ using the following loss function:
$\theta^{*}=\arg\min_{\theta}\ E_{t\sim u(\epsilon,1),\mathbf{x_{t}}\sim p(\mathbf{x_{t}}|\mathbf{x_{0}}),\mathbf{x_{0}}\sim\mathbf{p_{data}}}[||\mathbf{s_{\theta}}(\mathbf{x_{t}},t)-\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}}|\mathbf{x_{0}})||^{2}_{2}],$ (5)
@@
where network parameters, $\theta$, are estimated through a training process and the expected loss is evaluated over the uniform distribution timestep $t\in(0,T]$ and degraded images $\mathbf{x_{t}}$. The degraded images are generated by perturbing ground truth images $\mathbf{x_{0}}$ from a training dataset for the corresponding timestep $t$. Plugging the score estimate into Equation (4) yields the resulting reverse-time SDE:
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)+\nabla_{\mathbf{x_{t}}}\log p(\mathbf{y}|\mathbf{x_{t}}))]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (6)
@@
Following Chung et al.[13], the gradient of the likelihood term, $\nabla_{\mathbf{x_{t}}}\log p(\mathbf{y}|\mathbf{x_{t}})$, can be approximated as follows
@@
$\nabla_{\mathbf{x_{t}}}\log p(\mathbf{y}|\mathbf{x_{t}})=\nabla_{\mathbf{x_{t}}}\mathbf{\hat{x}_{0}}(\mathbf{x_{t}})\nabla_{\mathbf{x_{0}}}\log p(\mathbf{y}|\mathbf{x_{0}}=\mathbf{\mathbf{\hat{x}_{0}}})\;\;\;\text{where}\;\;\;\mathbf{\hat{x}_{0}}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)),$ (7)
@@
where $\mathbf{\hat{x}_{0}}$ denotes an approximation of $\mathbf{x_{0}}$ from $\mathbf{x_{t}}$, and $\nabla_{\mathbf{x_{t}}}$ denotes the gradient with respect to $\mathbf{x_{t}}$. For $\nabla_{\mathbf{x_{0}}}\log p(\mathbf{y}|\mathbf{x_{0}}=\mathbf{\hat{x}_{0}})$, we propose to apply a nonlinear physical model defined by a Gaussian distribution:
$\mathbf{y}\sim N(\mathbf{\bar{y}},\mathbf{D}\{\mathbf{y}\})\;\;\;\text{where}\;\;\;\mathbf{\bar{y}}(\mathbf{x_{0}})=\mathbf{B}\exp(-\mathbf{A}\mathbf{x_{0}}),$ (8)
where $\mathbf{\bar{\mathbf{y}}}$ denotes the mean measurement vector, $\mathbf{y}$ denotes the real projection coming from different CT system, $\mathbf{x_{0}}$ denotes the image of attenuation coefficients in transmission tomography, $\mathbf{A}$ is the linear projection matrix defined for a given CT imaging system, $\mathbf{D\{\cdot\}}$ represent a diagonal matrix. $\mathbf{B}$ is a matrix operator that can model various physical effects including measurement-dependent fluence levels, detector blur, etc. as in Tilley et al. [19] Given independent Gaussian noise with measurement variance $\mathbf{\sigma_{y}}=\mathbf{y}$ (approximating Poisson noise), one may write the measurement log-likelihood as
$\log p(\mathbf{y}|\mathbf{{x}_{0}})=-\frac{1}{2}(\mathbf{y}-\bar{\mathbf{y}}(\mathbf{{x}_{0}}))^{T}\mathbf{D}^{-1}\{\mathbf{y}\}(\mathbf{y}-\bar{\mathbf{y}}(\mathbf{{x}_{0}})),$ (9)
which includes the explicit dependence of the mean measurement model, $\bar{\mathbf{y}}$, on the underlying true image, $\mathbf{x_{0}}$, as defined in (8). Applying the chain rule, the gradient of this equation with respect to $\mathbf{x_{0}}$ is
$\nabla_{\mathbf{{x}_{0}}}\log p(\mathbf{y}|\mathbf{{x}_{0}})=\bar{\mathbf{y}}^{\prime}(\mathbf{{x}_{0}})^{T}\mathbf{D}^{-1}\{\mathbf{y}\}(\mathbf{y}-\bar{\mathbf{y}}(\mathbf{{x}_{0}})))$ (10)
where
$\bar{\mathbf{y}}^{\prime}(\mathbf{{x}_{0}})=-\mathbf{A}^{T}\mathbf{D}\{\exp(-\mathbf{A}\mathbf{{x}_{0}})\}\mathbf{B}^{T}.$ (11)
Combining terms and plugging in for the mean measurement model, we may write the expression
$\nabla_{\mathbf{x_{0}}}\log p(\mathbf{y}|\mathbf{x_{0}})=-\mathbf{A}^{\text{T}}\mathbf{D}\{\exp(-\mathbf{A}\mathbf{x_{0}})\}\mathbf{B}^{\text{T}}\mathbf{D}\{1/\mathbf{y}\}(\mathbf{y}-\mathbf{B}\exp(-\mathbf{A}\mathbf{x_{0}})),$ (12)
@@
where {}$\{\cdot\}^{\text{T}}$ denotes a matrix transpose with $\mathbf{A}^{\text{T}}$ representing backprojection. Equations (7) and (12) are plugged into Equation (4) to enable nonlinear physical model based diffusion posterior sampling in a reverse-time diffusion process.
2.2 Unconditional Training in Forward Diffusion
@@
A neural network is used to estimate the score function in the unconditional score-based diffusion model [10]. This network may be trained in an unsupervised fashion independent of the measurement model. The resulting DPS Nonlinear method does not require any extra training for different CT systems. In this work, we adopt a U-Net [20] with an attention residual module and time embedding module[10] for score estimation. The XCAT phantom[21] was used for the dataset of prior score model training. A total of 54 human models were divided into training, validation, and test sets with a ratio of approximately 8:1:1 and then individual slices of 256 $\times$ 256 voxels were extracted for each dataset. Individual human models were divided across datasets to avoid mixing between training, validation, and testing. For the stochastic process, we adopted a schedule defined by $\bar{\alpha}_{t}=1e^{-5t}$ with $t\in(0,T]$ where $T=1$. For training, we applied Adam optimizer, setting the batch size to 8 and the initial learning rate to 0.001. We trained our score estimator over 600 Epochs with early stopping to get the best weights for the prior score function. This trained score model was subsequently used for CT reconstruction for different CT system models.

### v2

# CT Reconstruction using Diffusion Posterior Sampling conditioned on a Nonlinear Measurement Model

@@ Abstract
Diffusion models have been demonstrated as powerful deep learning tools for image generation in CT reconstruction and restoration. Recently, diffusion posterior sampling, where a score-based diffusion prior is combined with a likelihood model, has been used to produce high quality CT images given low-quality measurements. This technique is attractive since it permits a one-time, unsupervised training of a CT prior; which can then be incorporated with an arbitrary data model. However, current methods rely on a linear model of x-ray CT physics to reconstruct or restore images. While it is common to linearize the transmission tomography reconstruction problem, this is an approximation to the true and inherently nonlinear forward model. We propose a new method that solves the inverse problem of nonlinear CT image reconstruction via diffusion posterior sampling. We implement a traditional unconditional diffusion model by training a prior score function estimator, and apply Bayes rule to combine this prior with a measurement likelihood score function derived from the nonlinear physical model to arrive at a posterior score function that can be used to sample the reverse-time diffusion process. This plug-and-play method allows incorporation of a diffusion-based prior with generalized nonlinear CT image reconstruction into multiple CT system designs with different forward models, without the need for any additional training. We develop the algorithm that performs this reconstruction, including an ordered-subsets variant for accelerated processing and demonstrate the technique in both fully sampled low dose data and sparse-view geometries using a single unsupervised training of the prior.
keywords:
CT reconstruction, deep learning reconstruction, diffusion model, diffusion posterior sampling, deep learning
[…]
@@ 1 Introduction
Neural networks have been widely investigated for CT reconstruction and have shown impressive performance capabilities [1, 2, 3]. Recently, diffusion models have been shown to be particularly powerful deep learning tools for CT reconstruction and restoration [4, 5, 6, 7, 8]. Such methods have generally been based on denoising diffusion probabilistic models (DDPM) [9] and score-based generative diffusion models through stochastic differential equations (SDEs) [10]. Most diffusion models for CT reconstruction have been trained in a supervised manner, where the conditional data input is known in both training and generation [7, 8, 11]. Recent research on image denoising and restoration has demonstrated the potential of unsupervised training by leveraging the power of posterior sampling, where the conditional data input is only available during the reverse-time diffusion process [12, 13, 5]. The unsupervised training does not assume a fixed measurement process during training, and can thus be flexibly incorporated with different measurement models without retraining. In these efforts, measurement model for diffusion posterior sampling has been presumed to be linear. Lopez-Montes et al.[14] have recently applied such an approach using a linearized model to tomographic data.
@@
Nevertheless, it is well-known that the transmission tomography forward model that relates the underlying attenuation values to the measurements is nonlinear. While the model is commonly linearized for mathematical convenience, such approximation will lose the opportunity to precisely describe those more sophisticated imaging system, thus can result in deviation from the true underlying data model and can cause problems in many circumstances including models for very low dose data [15], beam-hardening [16], and detector blur [17]. In this work, we propose an approach for CT image reduction using diffusion posterior sampling that is conditioned on a nonlinear measurement model (which we will refer to as DPS Nonlinear). Specifically, we seek to solve the inverse problem of CT image reconstruction via DPS with a score-based diffusion prior and a nonlinear physics-driven measurement likelihood term. This method guides the updates of the reverse-time diffusion process so that the final results are consistent with the raw (nonlinear) measurements from a CT imaging system. Our approach trains a traditional score-based diffusion model for unconditional CT image generation and applies Bayes rule by adding the gradient of the log-likelihood, i.e., from the non-linear physical model to achieve the posterior score function necessary for DPS. The method is plug-and-play, meaning it does not require any extra training for application to a different CT system measurement model. A preliminary development and evaluation of these proposed methods was presented in Li et al.[18] Significant changes from that presentation include training on a more realistic anatomy, adoption of a Poisson noise model, and more in-depth evaluations including comparisons to related deep learning approaches.
2 Method
2.1 Diffusion Posterior Sampling conditioned on a Nonlinear Physical Model
@@
Figure 1 shows the general workflow of our proposed DPS Nonlinear method, which combines diffusion posterior sampling with innovative likelihood updates based on a nonlinear physical measurement model (indicated in orange). This approach requires estimation of the prior distribution of CT images using a score-based generative model [9, 10]. Specifically, a diffusion model is adopted wherein a forward stochastic process (top portion of Figure 1) is defined where an image is perturbed by successively added noise. A reverse process (Figure 1, bottom) is then defined where we may sample images from the estimated prior distribution starting from a noise image.
Figure 1: The workflow of the proposed DPS Nonlinear method. The blue arrow represents a forward stochastic process defined by the equation in the top wherein $\mathbf{x_{0}}$, a “clean” ground truth CT image is successively degraded in intermediate noisy images, $\mathbf{x_{t}}$, indexed by time $t$, to an image of Gaussian noise, $\mathbf{x_{T}}$,. A reverse stochastic process is defined at the bottom that is conditioned on measurement data, $\mathbf{y}$. The orange arrow represents this conditioning and is based on a general nonlinear physical measurement model.
Specifically, we follow prior work [9, 10] where the forward diffusion process is governed by
$\mathbf{x_{t}}=\sqrt{\bar{\alpha}_{t}}\mathbf{x_{0}}+{\sqrt{(1-\bar{\alpha}_{t})}}\mathbf{\epsilon_{noise}}\;\;\;t\in(0,T],$ (1)
@@
where $\mathbf{\epsilon_{noise}}\sim\mathcal{N}(0,I)$ is Gaussian noise, $\mathbf{x_{0}}\sim\mathbf{p_{data}}$ denote a true image sampled from a training dataset, $\bar{\alpha}_{t}$ is a variance scheduler such that the final noise is close to a predefined noise distribution $\mathbf{\pi(x)}$, and the process is defined on $t\in(0,T]$. The final sample $\mathbf{p_{t=T}}$ usually follows the predefined noise distribution $\mathbf{\pi(x)}$ and it does not contain any information of training dataset $\mathbf{p_{data}}$ or $\mathbf{p_{t=0}}$. To recover the data in a tractable distribution, which can be achieved by reversing the perturbation forward process in Equation (1) to get the reverse-time SDE given by
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}})]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T],$ (2)
@@
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}}|\mathbf{y})]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (3)
@@
$\mathbf{d\mathbf{x_{t}}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}})+\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{t}}))]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (4)
@@
Following previous score matching methods[10, 19], the prior score $\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}})$ is estimated by a neural network called score model $\mathbf{s_{\theta}}(\mathbf{x_{t}},t)$ using the following loss function:
$\theta^{*}=\arg\min_{\theta}\ E_{t\sim u(\epsilon,1),\mathbf{x_{t}}\sim p(\mathbf{x_{t}}|\mathbf{x_{0}}),\mathbf{x_{0}}\sim\mathbf{p_{data}}}[||\mathbf{s_{\theta}}(\mathbf{x_{t}},t)-\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}}|\mathbf{x_{0}})||^{2}_{2}],$ (5)
@@
where the network parameters, $\theta$, are estimated through a training process and the expected loss is evaluated over the uniform distribution timestep $t\in(0,T]$ with degraded images $\mathbf{x_{t}}$. The degraded images are generated in the forward process in Equation (1) by perturbing ground truth images $\mathbf{x_{0}}$ from the training dataset for the corresponding timestep $t$. Plugging the score estimate into Equation (4) yields the resulting reverse-time SDE:
$\mathbf{dx_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)+\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{t}}))]dt+g(t)\mathbf{d\bar{w}_{t}}\;\;\;t\in(0,T].$ (6)
@@
Following Chung et al.[13], the gradient of the likelihood term, $\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{t}})$, can be approximated as follows:
@@
$\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{t}})=\nabla_{\mathbf{x_{t}}}\mathbf{\hat{x}_{0}}(\mathbf{x_{t}})\nabla_{\mathbf{x_{0}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{0}}=\mathbf{\mathbf{\hat{x}_{0}}})\;\text{where}\;\mathbf{\hat{x}_{0}}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)),$ (7)
@@
where $\mathbf{\hat{x}_{0}}$ denotes an approximation of $\mathbf{x_{0}}$ from $\mathbf{x_{t}}$, and $\nabla_{\mathbf{x_{t}}}$ denotes the gradient with respect to $\mathbf{x_{t}}$. For $\nabla_{\mathbf{x_{0}}}\mathrm{log}p(\mathbf{y}|\mathbf{x_{0}}=\mathbf{\hat{x}_{0}})$, we propose to apply a nonlinear physical model defined by a Poisson distribution:
${\mathbf{y}\sim Poisson\{\mathbf{\bar{y}}\}\;\;\;\text{where}\;\;\;\mathbf{\bar{y}}(\mathbf{x_{0}})=I_{0}\exp(-\mathbf{A}\mathbf{x_{0}}),}$ (8)
where $\mathbf{\bar{\mathbf{y}}}$ denotes the mean measurement vector, $\mathbf{y}$ denotes the real projection coming from different CT system, $\mathbf{x_{0}}$ denotes the image of attenuation coefficients in transmission tomography, $\mathbf{A}$ is the linear projection matrix defined for a given CT system geometry. $I_{0}$ is a scalar defining the mean number of photons/detector pixel in bare beam. Based on Erdogan et al.[20], we may write:
${\nabla_{\mathbf{x_{0}}}\mathrm{log}p(\mathbf{y}|\mathbf{{x}_{0}})=-\mathbf{A}^{T}\mathbf{y}+\mathbf{A}^{T}D\{I_{0}\exp(-\mathbf{A}\mathbf{{x}_{0}})\}}.$ (9)
@@
where {}$\{\cdot\}^{\text{T}}$ denotes a matrix transpose with $\mathbf{A}^{\text{T}}$ representing backprojection. Equations (7) and Equations (9) are plugged into Equation Equations (6) to enable nonlinear physical model based diffusion posterior sampling in a reverse-time diffusion process.
2.2 Training of Score Model in Forward Diffusion
@@
CT Lymph Nodes Dataset [21] was used for training of score model. A total of 156 subjects were divided into training, validation, and test sets with a ratio of approximately 7:2:1 and then individual slices of 512 $\times$ 512 voxels were then extracted for each dataset. We adopted a schedule defined by $\bar{\alpha}_{t}=1e^{-5t}$ with $t\in(0,T]$ where $T=1$. For the score model, we adopt a U-Net [22] with an attention residual module and time embedding module[10]. For training, we applied Adam optimizer, setting the batch size to 8 and the initial learning rate to 0.001. We trained our score model on training set over 1200 Epochs with early stopping of 100 Epochs. This trained score model was subsequently used for CT reconstruction for different CT system models.